### v2

# Ziwei Shi1,2     Xiaoran Zhang1,2     Yan Xia3,4     Yu Zang1,2     Siqi Shen1,2     Cheng Wang1,2

Abstract
We tackle the challenge of LiDAR-based place recognition, which traditionally depends on costly and time-consuming prior 3D maps. To overcome this, we first construct XA-L&RSI dataset, which encompasses approximately $110,000$ remote sensing submaps and $13,000$ LiDAR point cloud submaps captured in urban scenes, and propose a novel method, L2RSI, for cross-view LiDAR place recognition using high-resolution Remote Sensing Imagery. This approach enables large-scale localization capabilities at a reduced cost by leveraging readily available overhead images as map proxies. L2RSI addresses the dual challenges of cross-view and cross-modal place recognition by learning feature alignment between point cloud submaps and remote sensing submaps in the semantic domain. Additionally, we introduce a novel probability propagation method based on particle estimation to refine position predictions, effectively leveraging temporal and spatial information. This approach enables large-scale retrieval and cross-scene generalization without fine-tuning. Extensive experiments on XA-L&RSI demonstrate that, within a $100km^{2}$ retrieval range, L2RSI accurately localizes $83.27\%$ of point cloud submaps within a $30m$ radius for top-$1$ retrieved location. We provide a video to more vividly display the place recognition results of L2RSI at https://shizw695.github.io/L2RSI/.
Figure 1: (Left) We propose L2RSI for cross-view LiDAR-based place recognition without 3D pre-built map. Given a point cloud submap representing the surroundings of the robot and the absolute orientation provided by a LiDAR and a magnetometer, L2RSI provides the most probable location in a large-scale city using high-resolution remote sensing imagery. (Right) Place recognition performance in different retrieval ranges. Notably, L2RSI achieved a recall ($\!<\!30m$) of over $80\%$ at Top-1 retrieval within a range of $100km^{2}$.
1 Introduction
Place recognition aims to retrieve the closest match and its location from a pre-built database within the global coordinate system when GPS is weak or even denied, serving as an essential task for autonomous driving and robotic navigation. LiDAR-based place recognition is now increasingly attractive since 3D point clouds from LiDAR are invariant to lighting and shadows [12].
Conditions Tasks
GPS Range $100km^{2}$
IMU Database >$60000$
Query LiDAR+MAG Distance <$30m$
Reference RSI $Recall@1$ >$80\%$
Table 1: List of conditions and tasks.
Existing LiDAR-based place recognition methods face a significant limitation: they depend on up-to-date prior 3D maps, the acquisition and maintenance of which are time-consuming and costly. Recent research has sought to address this issue by exploring alternative approaches. Tang et al. [9, 10] unified overhead imagery and LiDAR into a pseudo point cloud modality, achieving retrieval-based place recognition. However, their approach only works on the assumption of a known route, which confines the retrieval process to a very limited range. Cho et al. [2] conducted LiDAR-based place recognition using OpenStreetMap (OSM). Due to the dual challenges of cross-view and cross-modality, existing methods typically restrict the retrieval area to a small range. The most commonly used 3D datasets include the Oxford Radar RobotCar [1], KITTI [4], and KITTI-360 [13]. The former consists of multiple repeated trajectories, covering an area of less than $2km^{2}$, while the latter two datasets are composed of several independent trajectories, with the largest covering range approximately $1km^{2}$.
As shown in Table 1, our task is to utilize high-resolution remote sensing imagery (RSI) as the database (due to the advantages of global coverage, cost-effectiveness, and timeliness), to achieve place recognition (LPR) in the range of over $100km^{2}$ urban scenes, more than $60000$ RSI and $30m$ required distance threshold, based on the only sensors of LiDAR and magnetometer (MAG).
In fact, using high-resolution remote sensing as the reference map will lead to a large gap between the query and databases in terms of domain and view, which will cause two major problems: (1) Most of the current retrieval solutions can hardly learn the correspondence between the query point cloud and the remote sensing image retrieval database directly. (2) Single query can be ambiguous and unstable, which limit the practicality of the system, not to mention the retrieving range of over $100km^{2}$.
To address the first problem, we observe that the content and style between LiDAR point clouds and remote sensing imagery, even from the same location, can differ significantly. However, they exhibit strong correlations in the semantic domain, as illustrated in the bottom-left corner of Fig. 1. Inspired by this observation, we propose a framework, named L2RSI, which includes a semantic contrastive learning network to align the global descriptors of LiDAR point cloud submaps and remote sensing submaps in the semantic space. The point cloud submaps are obtained through the registration of short-range LiDAR scans, overcoming the sparsity of single-frame point clouds. With the direction provided by an additional magnetometer (MAG), these submaps are transformed into a bird’s-eye view (BEV) representation through rotation and projection, enabling the network to effectively exploit the shared semantic features visible in both modalities.
Specific to the second problem, we propose the Spatial-Temproal Particle Estimation (STPE) to aggregate spatio-temporal information of continuous queries, thus to eliminate the ambiguity between geometrically distant but semantically similar positions. Specifically, the retrieval results in each query are considered as particles. Instead of filtering out the incorrect particles, our idea is to aggregate these particles and estimate a more accurate probability density distribution for the current position. Distribution of the particles of a single query can be modeled as the mixture of multiple gaussian models (GMMs). Then, the aggregation spread over a desired time window through an estimated inter-frame correspondence of the point clouds, estimating the probability density of the current query and refining the retrieval results.
Our contributions are summarized as follows:
  • To the best of our knowledge, the proposed framework is the first to address the challenge of large-scale (over $100km^{2}$) urban cross-view LiDAR-based place recognition with high-resolution remote sensing imagery, demonstrating promising performance with significant practical applicability.
  • We propose a particle estimation algorithm that utilizes the mixture of multiple gaussian models to aggregate spatio-temporal information and infer the probability density of the current position, significantly improving the retrieval performance.
  • The proposed method exhibits remarkable generalization ability, achieves $Recall@1$ ($\!<\!30m$) of $83.27\%$ within a retrieval range of $100km^{2}$, and demonstrates promising potential on cross scene generalization without fine-tuning.
2 Problem statement
We begin by defining the large-scale remote sensing imagery $M_{ref}\!=\!\{m_{i}\}_{i=1}^{M}$ to be a collection of square submaps $m_{i}$. Each submap is sampled at equal intervals with partial overlap and includes the UTM coordinates of the center point. Let $Q$ be a query point cloud submap obtained by aggregating a short sequence of LiDAR scans. Our task is defined as determining the coarse position of the center point in the query point cloud submap, which is a cross-view, cross-modal retrieval problem:
$m^{*}=\underset{m_{i}\in M_{ref}}{argmin}\,Sim(F_{q}(Q,o),F_{ref}(m_{i})),$ (1)
where $F_{q}(\cdot)$ and $F_{ref}(\cdot)$ denote the deep learning networks used to encode the point cloud submap $Q$ and remote sensing submaps $m_{i}$ into a shared feature space. $Sim(\cdot,\cdot)$ represents a measure of similarity, e.g., the dot product. $o$ represents the orientation of the intermediate frame in $Q$. We obtain a noisy measurement $o$ from a magnetometer.
3 Methodology
Fig. LABEL: shows the overview of the proposed L2RSI. We will introduce the Data preprocessing module in Sec. LABEL:. Sec. LABEL: givess the approach to train the global descriptor extractor for the query point cloud submaps and the remote sensing submaps. Sec. LABEL: introduces STPE, which aims to estimate the probability density of the current query position and refine these candidates.
4 Experiment
4.1 Benchmark dataset
(a) XA-L&RSI
(b) Oxford-L&RSI
Figure 2: Dataset configuration. In XA-L&RSI 2(a), different colored rectangles highlight test sets from databases of varying sizes, all sharing the same query trajectory (green). Besides, the training trajectory (light blue) and discarded trajectory at the boundary (red) are annotated. In Oxford-L&RSI 2(b), the green trajectory is as the query for the test set, while the entire remote sensing imagery serves as the database.
Due to the lack of publicly available datasets that provide correspondences between large-scale outdoor LiDAR scans and remote sensing imagery, we construct the XA-L&RSI and Oxford-L&RSI datasets for training and evaluation.
XA-L&RSI. We purchase high-resolution remote sensing imagery covering Xiang’an District from AIRSAT Technology Group, with a resolution of $0.5m$. We extract submaps of $60m\!\times\!60m$ from the original remote sensing imagery at intervals of $40m$, resulting in a one-third overlap between adjacent submaps. And the submaps lacking road semantics in their $30m\!\times\!30m$ central region are filtered out.
We utilize a Ouster OS1-64 LiDAR to acquire point cloud data in the Xiang’an District of Xiamen, covering a trajectory of approximately $100km$. We record directional information through a magnetometer with an average error of less than $10^{\circ}$. During point cloud preprocessing, we construct a point cloud submap every $5m$ and compress it into a semantic BEV image with a side length of $d=60m$. There is a three-year temporal gap between LiDAR data and remote sensing imagery.
The vehicle trajectory is shown in Fig. 2(a). We select approximately $5km$ of the trajectory as the test set, with the remaining portion used for training. For the training set, we manually exclude data with inherent misalignment between ground and aerial semantic information to avoid confusion during training. This includes the areas where the sky is obstructed, such as the road segments beneath overpasses and dense street trees, as well as changes in buildings, vegetation, and roads due to the temporal gap. For the test set, we established databases of $4km^{2}$ (S), $9km^{2}$ (M), $16km^{2}$ (L) and $100km^{2}$ (G) to evaluate the retrieval performance of the proposed method, see Table 2. Notably, the point cloud at the boundary between the test and training sets are discarded, ensuring that the network is never exposed to the test set data during training.
XA-L&RSI Submap number Area ($km^{2}$)
PC RSI
Training Set 12194 47913 65
Test Set-S 862 3094 4
Test Set-M 862 6479 9
Test Set-L 862 12119 16
Test Set-G 862 63047 100
Oxford-L&RSI Submap number Area ($km^{2}$)
PC RSI
11-14-02-26 1697 2103 2.8
14-12-05-52 1671 2103 2.8
Table 2: Summary statistics for dataset.
[…]
Test Set-S Test Set-M Test Set-L Test Set-G 11-14-02-26 14-12-05-52
GeoDTR [15] 8.68/25.27 6.25/19.36 4.74/14.95 4.39/9.26 0.47/3.18 0.48/3.71
Sample4Geo [3] 28.63/53.92 25.39/49.75 24.23/46.27 22.95/43.02 7.13/20.68 7.51/21.75
FRGeo [14] 22.26/49.88 19.72/42.45 17.16/38.05 15.54/34.45 1.62/7.34 1.58/7.86
L2RSI (w/o STPE) 30.05/68.56 23.90/61.95 21.93/57.42 22.62/60.32 10.19/30.23 11.25/30.58
L2RSI (w. SuperGlobal [8]) 30.39/57.54 23.90/53.02 22.27/48.26 20.19/42.92 8.9022.98 8.98/22.86
L2RSI (w. PF) 71.66/86.88 55.17/72.47 50.87/72.47 47.85/58.65 21.40/31.43 16.83/29.46
L2RSI (w. STPE) 88.93/92.13 87.95/91.64 85.49/90.65 83.27/87.82 42.29/59.41 43.77/62.76
[…]
Oxford-L&RSI. Oxford Radar RobotCar is a famous urban scene dataset [1] for localization. The data is collected with a Velodyne HDL-32E LiDAR. The dataset includes multiple repeated traversals, each covering about $10km$, as shown in Fig. 2(b). Based on this, we supplement the database with the corresponding remote sensing imagery, covering an area of approximately $2.8km^{2}$. And we simulate the magnetometer by adding random noise within ±$\pm 15^{\circ}$ to the ground truth of orientation, as the official dataset does not include a magnetometer. The source and resolution of remote sensing imagery, as well as the data preprocessing process, are the same as those of the XA-L&RSI. The data of 11-14-02-26, 14-12-05-52 are used as the test set to evaluate the generalization ability of the proposed method without fine-tuning, see Table 2.
4.2 Evaluation criteria
Following pervious place recognition task [11, 6], we use retrieval recall to evaluate LiDAR-based place recognition based on remote sensing imagery. $Recall@N$ means the recall at top-$N$ retrieval results, that is, the proportion of queries for which the correct result is retrieved within the top-$N$ results. We define a correct result as when the Euclidean distance between the center of the retrieved remote sensing submap and the center of the query point cloud submap is less than $30m$. We particularly emphasize on $Recall@1$, as it has the most direct practical relevance.
4.3 Implementation details
The proposed L2RSI is implemented with Pytorch [7]. The network is trained with Adam optimizer [5]. All experiments are performed on Ubuntu $20.04.1$ with a single NVIDIA GTX $3090$ GPUs with 24GB of memory.
4.4 Results
4.4.1 Evaluation on test sets
The proposed L2RSI is the first to perform cross-view LiDAR-based place recognition in large-scale urban scenes with high-resolution remote sensing imagery. Several state-of-the-art cross-view image retrieval networks are capable of handling preprocessed cross-modal semantic images, e.g. GeoDTR [15], Sample4Geo [3] and FRGeo [14]. Firstly, we benchmark them to evaluate the performance of our method in the test sets of XA-L&RSI with different size databases. The training settings on XA-L&RSI for all methods are consistent, except for the image size and descriptor dimensions, which follow the official configurations. The comparison is summarized in Table 3. L2RSI (w. STPE) significantly outperforms the baseline methods, reaches a $Recall@1$ of $88.93\%$ with a database of $4km^{2}$. As the retrieval range expands, L2RSI (w. STPE) demonstrates increased robustness, with the database area increased to $100km^{2}$ and only a $5.66\%$ degradation in $Recall@1$ and a $4.31\%$ degradation in $Recall@10$. In addition, we also introduced a comparison of two refinement strategies, including SuperGlobal reranking [8] and classical particle filtering (PF). The former focuses on refining individual queries and database features, leading to only marginal improvements in performance. For the latter, we apply its principles within the retrieval framework, initializing the number of particles to $120$. With the integration of sequence information, the performance of L2RSI (w. PF) shows considerable improvement. However, in challenging cross-view and cross-modal retrieval tasks, particles frequently decay, necessitating re-initialization and impeding the flow of sequence information. The proposed STPE substitutes filtering with aggregation, efficiently utilizing spatio-temporal information to substantially enhance the unstable performance of single query.
Secondly, we directly test the model on the Oxford-L&RSI without any fine-tuning. Given the stark differences in equipment models, road widths, architectural styles, etc., this challenge marks the highest level of generalization ability. In the lower part of Table 3, the baseline methods fail completely, while L2RSI (STPE) achieves $Recall@1$ over $40\%$ and $Recall@10$ above $60\%$ on both trajectories. Although the current results are not yet sufficient for real-world deployment, they demonstrate great potential for future applications.
5 Conclusion
We present L2RSI, the first method for cross-view LiDAR-based place recognition within large-scale (over $100km^{2}$) urban scenes using high-resolution remote sensing imagery. We unify LiDAR point cloud submaps and remote sensing submaps into a shared semantic space via a carefully-designed contrastive learning network. Furthermore, we propose the Spatial-Temproal Particle Estimation to aggregate the spatio-temporal constraints of sequential queries, thereby further refining the performance of global place recognition. Extensive experiments on the XA-L&RSI and Oxford-L&RSI datasets demonstrate the effectiveness and generalization ability of L2RSI. Future work will focus on reducing the reliance on additional magnetometers and further addressing the challenges of deploying this technology in real-world applications.